Discuss Poynting vector related to power flow in free space. A uniform plane wave at frequency of 200 MHz travel in vacuum along positive Y direction. The electric field at some instance is given as  . Find the phase constant of the wave.


1
Expert's answer
2021-04-19T17:39:36-0400

To be given in question

Frequency n=200MHz

Wave move y direction

To be asked in question

Phase constant of wave = ?

We know that

Plane wave equation

"E=E_{0}sin(wt-ky) j"

"(wt-ky)=\phi =phase angle"

"(2\pi\times 200\times10^6t -ky)=\phi"

"(4\pi\times10^8t-ky)=\phi" "\rightarrow(1)"

equation (1) is called phase constant

Phase constant varring with respect time